# Predicting O-GlcNAcylation Sites in Mammalian Proteins with Transformers and RNNs Trained with a New Loss Function

### v1

@@ Abstract
Glycosylation, a protein modification, has multiple essential functional and structural roles. O-GlcNAcylation, a subtype of glycosylation, has the potential to be an important target for therapeutics, but methods to reliably predict O-GlcNAcylation sites had not been available until 2023; a 2021 review correctly noted that published models were insufficient and failed to generalize. Moreover, many are no longer usable. In 2023, a considerably better RNN model with an F score of 36.17% and an MCC of 34.57% on a large dataset was published. This article first sought to improve these metrics using transformer encoders. While transformers displayed high performance on this dataset, their performance was inferior to that of the previously published RNN. We then created a new loss function, which we call the weighted focal differentiable MCC, to improve the performance of classification models. RNN models trained with this new function display superior performance to models trained using the weighted cross-entropy loss; this new function can also be used to fine-tune trained models. A two-cell RNN trained with this loss achieves state-of-the-art performance in O-GlcNAcylation site prediction with an F score of 38.82% and an MCC of 38.21% on that large dataset.
@@ 1 Introduction
Glycosylation, a co- and post-translational modification, occurs when glycan(s) are added to proteins. Glycosylation is important functionally and structurally (Schjoldager et al., 2020; Chang et al., 2020); conversely, incorrect glycosylation or deglycosylation is associated with multiple diseases such as cancers (Stowell et al., 2015), infections (Bhat et al., 2019), and congenital disorders (Jaeken, 2013). O-linked glycosylation occurs when a glycan is added to an oxygen of an amino acid (usually serine or threonine in mammals). O-GlcNAcylation, a subtype of O-linked glycosylation, occurs when the first glycan added is an N-Acetylglucosamine (abbreviated GlcNAc) (Schjoldager et al., 2020). O-GlcNAcylation is catalyzed solely by the enzymes OGT and OGA.
The potential for diagnoses and treatments of glycosylation is very high, potentially benefiting the biomedical and pharmaceutical industry, physicians, and patients. For example, multiple carcinoma types have elevations in fucosylation, branching, and sialyation (Almeida & Kolarich, 2016). Disialoganglioside is expressed by the vast majority of neuroblastomas, and Phase I–III studies have shown that neuroblastoma may be treated by anti-disialoganglioside monoclonal antibodies (Ho et al., 2016; Ahmed & Cheung, 2014). Poly-2,8-sialyation can increase the half-lives of antibodies and does not cause tolerance problems (Van Landuyt et al., 2019). Contrarily, glycans not produced by humans, such as the CHO-cell-derived N-glycolylneuraminic acid (Padler-Karavani et al., 2008; Hokke et al., 1995), may hinder biotherapeutics (Hokke et al., 1995). According to recent research, O-GlcNAcylation can be a powerful target for therapeutics (Zhu & Hart, 2021), further emphasizing the relevance of glycosylation.
While glycosylation is very important for biotherapeutics, challenges still exist. Glycosylation investigations must consider both the position of and the glycan composition at each glycosylation site (Almeida & Kolarich, 2016). Analysis and elucidation of specific glycosylation functions are challenging due to the large variety of glycans and glycosylation sites (Schjoldager et al., 2020). The elaborate equipment used to characterize glycosylation limits clinical laboratories from analyzing patient samples on a large scale, hindering personalized medicine (Almeida & Kolarich, 2016).
@@
Glycosylation research can be assisted by computational tools, increasing biological understanding and allowing the prediction of glycosylation. In the context of glycosylation, classifier models (such as YinOYang (YoY) (Gupta & Brunak, 2002) or O-GlcNAcPRED-II (Jia et al., 2018)) may predict the location of N- or O-glycosylation sites. In that same context, regressor models (such as the models in Moon et al. (2021) and Seber & Braatz (2023a)) may quantitatively predict protein glycan distributions based on some input (such as enzyme levels). Previously, O-GlcNAcylation site prediction models had insufficient performance to help advance research in the area. Specifically, Mauri et al. (2021) found that no published model until then could achieve a precision 9% on a medium-sized independent dataset, suggesting that models to predict the location of O-GlcNAcylation sites failed to generalize successfully in spite of the high metrics achieved with their training data. The models evaluated in Mauri et al. (2021) also have low F scores and Matthew Correlation Coefficients (MCCs), another sign that their performance is insufficient. Seber & Braatz (2023b) used RNN models (specifically, LSTMs) to predict the presence of O-GlcNAc sites from mammalian protein sequence data. That work used two different datasets (from Mauri et al. (2021) and Wulff-Fuentes et al. (2021)), the latter of which is among the largest O-GlcNAcylation datasets available. These RNNs obtained vastly superior performance to the previously published models, achieving an F score more than 3.5-fold higher and an MCC more than 4.5-fold higher than the previous state of the art. Furthermore, Shapley values were used to interpret the predictions of that RNN model through the sum of simple linear coefficients (Shapley, 1951). These predictions with Shapley values maintained most of the performance of the original RNN models (Supplemental data of Seber & Braatz (2023b)).
@@
While this was a significant improvement over previous models and brought interpretability for the first time in a publication on data-driven O-GlcNAcylation site prediction models, the test-set F score and MCC achieved by the best RNN of Seber & Braatz (2023b) was 36.17% and 34.57%. Thus, there is room for improvement, and we believed the transformer architecture could lead to a better model, but this hypothesis proved false, as LSTMs outperform transformers. This work also develops and explores using a novel loss function that aims to directly improve the models’ MCCs, and compares the superior results attained with new loss function against results obtained with the weighted cross-entropy loss function. Rigorous 5-fold cross-validation is employed in the model training, ensuring the reliability of reported metrics and predictions. The improved model generated by this study is provided as open-source software, allowing the reproducibility of the work, the retraining of the models as additional or higher quality O-GlcNAcylation data become available, and the use of the models to further improve the understanding and applications of O-GlcNAcylation.
@@ 2 Materials and Methods
2.1 Datasets
The data used to train the model was first generated by Wulff-Fuentes et al. (2021) and was modified by Seber & Braatz (2023b). The modified version is available at github.com/PedroSeber/O-GlcNAcylation_Prediction/blob/master/OVSlab_allSpecies_O-GlcNAcome_PS.csv. This dataset contains 558,168 unique S/T sites from mammalian proteins. Out of these, 13,637 (2.44%) are O-GlcNAcylated. The same procedure as in Seber & Braatz (2023b) was employed to filter homologous and isoform proteins, a sequence selection based on a window size of 5 AA on each side of the central S/T (11 AA total), even for the larger windows. 20% of each dataset is separated for testing, with the remaining 80% used for cross-validation with five folds. This procedure follows what was previously done in Seber & Braatz (2023b).
@@ 2.2 Transformer Models
Transformer encoder models are constructed using PyTorch (Paszke et al., 2019) and other Python packages (Harris et al., 2020; Pedregosa et al., 2011; McKinney, 2010). Embed sizes (ES) {24, 60, or 120}, learning rates {1 or 1}, {2, 4, 6, or 8} stacked encoder cells, cell hidden sizes (widths) {ES/4, ES/2, ES, ES2, ES3, ES4, ES5}, {2, 3, 4, 5, or 6} attention heads111When available in a given embed size., post-encoder MLP sizes {0, ES/4, ES/2, ES}, and window sizes of {5, 10, 15, or 20} were used. Training is done in 100 epochs with the AdamW optimizer with a weight decay parameter {0, 1, 5, or 1} and cosine scheduling (Loshchilov & Hutter, 2017). Cosine scheduling has been used mostly in computer vision settings and has achieved significant results with imbalanced datasets (Kukleva et al., 2023; Mishra et al., 2019). The best combination of hyperparameters is determined by a semi-grid search222Combinations of hyperparameters that were a priori clearly inferior were not tested., and the combination with the highest cross-validation average F score is selected. Finally, the performance on an independent test dataset of a model using that combination of hyperparameters is reported. Multiple thresholds are used to construct a Precision-Recall (P-R) curve, and the F scores and MCCs of the models are analyzed, allowing for a throughout evaluation of the best model’s performance.
@@ 2.3 RNN Models
RNN models (specifically, LSTMs) are constructed using the same packages mentioned in Section 2.2. RNN hidden sizes {225 to 1575 in multiples of 75}, learning rates {1, 5, 1}, post-LSTM MLP sizes {37 and 75 to 825 in multiples of 75}, and window sizes of {20} were used. Training is done in 70 epochs with the AdamW optimizer with a weight decay parameter {0, 1, or 2} and cosine scheduling (Loshchilov & Hutter, 2017). Hyperparameter selection and testing / evaluation procedures are as described in Section 2.2.
@@ 2.4 Loss Functions
The cross-entropy (CE) loss is the most common loss function for classification tasks, and its weighted variant allows it to be used with unbalanced data. Despite its widespread use and efficacy, the CE loss has some important limitations. First, the CE loss is distinct from the evaluation metrics used. Users of classification models may desire to maximize the F score or MCC metrics, but the model with the lowest CE loss is not necessarily the one with the highest metrics. Moreover, a model can lower its CE loss by manipulating the predictions of examples that are already correctly separated. For an extreme example, consider a model that predicts 0.8 for all positive examples and 0.2 for all negative examples and a second model that predicts 0.99 for all positive examples and 0.01 for all negative examples. Both models perform perfect classification, yet the second model will have a lower CE loss. In a less extreme scenario, a model may improve its CE loss by simply manipulating the probabilities of easy-to-classify samples, but such manipulations would not improve metrics such as F score or MCC.
@@
There are multiple ways to remedy these issues. Lin et al. (2017) created the Focal Loss, which provides a correction factor that reduces the effect of increasing the certainty of a correct prediction, forcing the model to improve on difficult-to-classify samples to lower the focal loss. Berman et al. (2018) created the Lovász loss, which is an optimizable form of the Jaccardi Index and is claimed to perform better than the CE loss due to its better categorizing of small objects and lower false negatives. Other losses not evaluated in this work include the Baikal loss of Gonzalez & Miikkulainen (2020), the soft cross entropy of Ilievski & Feng (2017), the sigmoid F of Bénédict et al. (2022), the work of Lee et al. (2021), and the MCC loss of Abhishek & Hamarneh (2021). In particular, the approaches of this work resemble those of Lee et al. (2021) and Abhishek & Hamarneh (2021), but these references did not use weighting or focal transforms, and the definitions for the differentiable F and MCC loss functions used in these works are slightly different.
@@
The standard formulations of evaluation metrics (such as the F score or MCC) are not differentiable, so they cannot be used as loss functions. However, it is possible to modify them for differentiability. One method is treating the true / false categories as prediction probabilities instead of binary values. For example, if the model predicts for a positive (negative) sample, that sample would increase the true positives (negatives) by 0.7 and the false negatives (positives) by 0.3. This modification can be considered as a differentiable MCC loss function. The focal modification of Lin et al. (2017) can be added to this loss function by modifying to , which leads to the focal differentiable MCC loss function. Finally, it is possible to add class weighting: by multiplying the true positives and false negatives by a scalar W, it is possible to prioritize predictions on the positive class (W > 1) or negative class (W < 1). All the above changes combined lead to the weighted focal differentiable MCC loss function (Algorithm 1).
@@ 3.1 Transformer Encoders Have Good Predictive Power on This Dataset, but are Inferior to RNNs
The primary goal of this subsection is to compare transformer models with the best RNN from Seber & Braatz (2023b) using the modified dataset of Wulff-Fuentes et al. (2021) (Section 2.1) and the weighted CE loss function. While our transformer models surpass models published before Seber & Braatz (2023b) and achieve an F score equal to 24.31% and an MCC equal to 22.49% (Table 1), transformers exhibit inferior recall at the same precision level (Fig. 1) than the RNNs from Seber & Braatz (2023b), leading to inferior F and MCC metrics. The precision for our transformer models increases monotonically with increasing threshold, while the F score peaks at different thresholds for each model.
Metric Transformer-15; 5 win (This Work) Transformer-15; 10 win (This Work) Transformer-15; 15 win (This Work) Best Threshold 0.60 0.20 0.50 Recall (%) 20.54 26.06 26.02 Precision (%) 15.63 13.77 21.28 F Score (%) 17.75 18.02 23.41 MCC (%) 15.50 16.08 21.36
A surprising result is how the performance of the transformer models does not change significantly with varying hyperparameters. Using large hidden sizes or adding another linear layer after the encoders slightly degrades the performance, while increasing the number of attention heads (to 4–5) and the number of stacked encoder layers (to 4) marginally improves the model. The most surprising result is the lack of significant improvement with increasing window size. As in Seber & Braatz (2023b), our models perform better with larger window sizes for sizes up to 20 (Fig. 1 and Table 1). The performance of a transformer encoder with a window size = 5 is only marginally inferior to the RNN from Seber & Braatz (2023b) using the same window size. However, whereas the RNNs’ performances increase quite significantly with increasing window size, the performance of the transformers barely improves. The only significant performance increase comes from increasing the window size from 10 to 15, but barely any performance is gained from further increasing this window size to 20 (Fig. 1).

### v2

@@ Abstract
Glycosylation, a protein modification, has multiple essential functional and structural roles. O-GlcNAcylation, a subtype of glycosylation, has the potential to be an important target for therapeutics, but methods to reliably predict O-GlcNAcylation sites had not been available until 2023; a 2021 review correctly noted that published models were insufficient and failed to generalize. Moreover, many are no longer usable. In 2023, a considerably better recurrent neural network (RNN) model with an F1 score of 36.17% and a Matthews Correlation Coefficient (MCC) of 34.57% on a large dataset was published. This article first sought to improve these metrics using transformer encoders. While transformers displayed high performance on this dataset, their performance was inferior to that of the previously published RNN. We then created a new loss function, which we call the weighted focal differentiable MCC, to improve the performance of classification models. RNN models trained with this new function display superior performance to models trained using the weighted cross-entropy loss; this new function can also be used to fine-tune trained models. A two-cell RNN trained with this loss achieves state-of-the-art performance in O-GlcNAcylation site prediction with an F1 score of 38.88% and an MCC of 38.20% on that same large dataset.
@@ 1 Introduction
Glycosylation, a co- and post-translational modification, occurs when glycan(s) are added to proteins. O-linked glycosylation occurs when a glycan is added to an oxygen of an amino acid (usually serine or threonine in mammals). O-GlcNAcylation, a subtype of O-linked glycosylation, occurs when the first glycan added is an N-Acetylglucosamine (abbreviated GlcNAc) (Schjoldager et al., 2020). Unlike other forms of glycosylation, O-GlcNAcylation may be viewed similarly to phosphorylation, for no other glycans are further added to a site after GlcNAc and the process is catalyzed solely by two enzymes, OGT and OGA. O-GlcNAcylation is important functionally and structurally (Schjoldager et al., 2020; Chang et al., 2020); conversely, incorrect O-GlcNAcylation or its improper removal is associated with multiple diseases such as cancers (Shi et al., 2022), infections (Chang et al., 2020), and heart failure (Umapathi et al., 2021). According to recent research, O-GlcNAcylation can be a powerful target for therapeutics (Zhu & Hart, 2021), further emphasizing its relevance.
@@
While O-GlcNAcylation is essential for human health and very important for biotherapeutics, challenges still exist. Because O-GlcNAcylation is complete after the addition of a single glycan, the problem of predicting glycan distributions in a glycosylation site does not exist; thus, the main predictive task is to determine where and when (if ever) an amino acid will be O-GlcNAcylated. As summarized by Seber & Braatz (2024), this task is challenging for “multiple reasons”, which include “a low frequency of events (only about 2% of S/T sites are O-GlcNAcylated) and a lack of a motif to guide predictive efforts”. Moreover, Seber & Braatz (2024) also note that “the effects of neighboring amino acids likely influence whether an S/T is O-GlcNAcylated”, a fact corroborated by their interpretability studies done with their final model, but many previous works use model architectures that cannot directly take into account any type of sequential information. The pioneer classifier machine-learning model for this challenging task is YinOYang (YoY) (Gupta & Brunak, 2002), which also had one of the highest performances until the work of Seber & Braatz (2024), first published in 2023. However, until at least 2021, O-GlcNAcylation site prediction models had insufficient performance to help advance research in the area (Mauri et al., 2021). Specifically, Mauri et al. (2021) found that no published model until then could achieve a precision 9% on a medium-sized independent dataset, suggesting that models to predict the location of O-GlcNAcylation sites failed to generalize successfully in spite of the high metrics achieved with their training data. The models evaluated by Mauri et al. (2021) also have low F1 scores and Matthew Correlation Coefficients (MCCs), another sign that their performance is insufficient. Seber & Braatz (2024) used recurrent neural network (RNN) models (specifically, long short-term memory [LSTM] models) to predict the presence of O-GlcNAc sites from mammalian protein sequence data. That work used two different datasets (from Mauri et al. (2021); Wulff-Fuentes et al. (2021)), the latter of which is among the largest O-GlcNAcylation datasets available. These RNNs obtained vastly superior performance to the previously published models, achieving an F1 score more than 3.5-fold higher and an MCC more than 4.5-fold higher than the previous state of the art. Furthermore, Shapley values were used to interpret the predictions of that RNN model through the sum of simple linear coefficients (Shapley, 1951). These predictions with Shapley values maintained most of the performance of the original RNN models (Supplemental data of Seber & Braatz (2024)). Conversely, two other models that were also released recently, LM-OGlcNAc-Site (Pokharel et al., 2023) and O-GlcNAcPRED-DL (Hu et al., 2024), failed to obtain high metrics in independent test sets despite their complex architectures (an ensemble of transformers pre-trained for proteins in LM-OGlcNAc-Site’s case, and a CNN+LSTM model in O-GlcNAcPRED-DL’s case) and the direct use of some structural information. These two models still cannot reach precision 9% on independent test sets, and an analysis by Seber & Braatz (2024) (Table 4 of that work) noted that their model had a 2-fold performance improvement over O-GlcNAcPRED-DL, despite the fact that the latter was operating with a few advantages.
@@
While the work of Seber & Braatz (2024) was a significant improvement over previous models and brought interpretability for the first time in a publication on data-driven O-GlcNAcylation site prediction models, the test-set F1 score and MCC achieved by the best RNN of Seber & Braatz (2024) was 36.17% and 34.57%. Thus, there is room for improvement, and we believed the transformer architecture could lead to a better model, but this hypothesis proved false, as LSTMs outperform transformers. This work also develops and explores using a novel loss function that aims to directly improve the models’ MCCs, and compares the superior results attained with new loss function against results obtained with the weighted cross-entropy loss function. 5-fold cross-validation is employed in the model training, ensuring the reliability of reported metrics and predictions. The improved model generated by this study is provided as open-source software, allowing the reproducibility of the work, the retraining of the models as additional or higher quality O-GlcNAcylation data become available, and the use of the models to further improve the understanding and applications of O-GlcNAcylation.
@@ 2 Materials and Methods
2.1 Dataset
The data used to train the model was first generated by Wulff-Fuentes et al. (2021) and was modified by Seber & Braatz (2024). The modified version is available at github.com/PedroSeber/O-GlcNAcylation_Prediction/blob/master/OVSlab_allSpecies_O-GlcNAcome_PS.csv. This dataset contains 558,168 unique S/T sites from mammalian proteins. Out of these, 13,637 (2.44%) are O-GlcNAcylated. The same procedure as in Seber & Braatz (2024) was employed to filter homologous and isoform proteins, a sequence selection based on a window size of 5 AA on each side of the central S/T (11 AA total), even for the larger windows. 20% of each dataset is separated for testing, with the remaining 80% used for cross-validation with five folds. This procedure follows what was previously done in Seber & Braatz (2024). To ensure that the choice of train/cross-validation and test sets is not biased and to highlight the robustness of our architecture and novel loss function, this work also performs nested validation with the best model, similarly to Seber & Braatz (2024). In each nested validation round, 20% of the data are used as a test set and the remaining is used for hyperparameter selection via 5-fold cross-validation (as above).
@@ 2.2 Transformer Models
Transformer encoder models are constructed using PyTorch (Paszke et al., 2019) and other Python packages (Harris et al., 2020; Pedregosa et al., 2011; McKinney, 2010). Embed sizes (ES) {24, 60, or 120}, learning rates {1 or 1}, {2, 4, 6, or 8} stacked encoder cells, cell hidden sizes (widths) {ES/4, ES/2, ES, ES2, ES3, ES4, ES5}, {2, 3, 4, 5, or 6} attention heads (when available in a given ES), post-encoder MLP sizes {0, ES/4, ES/2, ES}, and window sizes of {5, 10, 15, or 20} were used. Training is done in 100 epochs with the AdamW optimizer with a weight decay parameter {0, 1, 5, or 1}, Kaiming uniform initialization (He et al., 2015) with , and cosine scheduling (Loshchilov & Hutter, 2017). Cosine scheduling has been used mostly in computer vision settings and has achieved significant results with imbalanced datasets (Kukleva et al., 2023; Mishra et al., 2019), and it was also used by Seber & Braatz (2024). The best combination of hyperparameters is determined by a semi-grid search111Combinations of hyperparameters that were a priori clearly inferior were not tested., and the combination with the highest cross-validation average F1 score is selected. Finally, the performance on an independent test dataset of a model using that combination of hyperparameters is reported. Multiple thresholds are used to construct a Precision-Recall (P-R) curve, and the F1 scores and MCCs of the models are analyzed, allowing for a throughout evaluation of the best model’s performance.
@@ 2.3 RNN Models
RNN models (specifically, LSTMs) are constructed using the same packages mentioned in Section 2.2. RNN hidden sizes {225 to 1575 in multiples of 75} for the first cell, {0 to 300 in multiples of 75} for the second cell, and {0 to 225 in multiples of 75} for the third and fourth cells, learning rates {1, 5, 1}, post-LSTM MLP sizes {37 and 75 to 825 in multiples of 75}, and window sizes of {20} were used. Training is done in 70 epochs with the AdamW optimizer with a weight decay parameter {0, 1, or 2}, Kaiming uniform initialization (He et al., 2015) with , and cosine scheduling (Loshchilov & Hutter, 2017). Hyperparameter selection and testing / evaluation procedures are as in Section 2.2.
@@ 2.4 Loss Functions
The cross-entropy (CE) loss is the most common loss function for classification tasks, and its weighted variant is suitable for unbalanced data. Despite its widespread use and efficacy, the CE loss has some important limitations. First, the CE loss is distinct from the evaluation metrics used. Users of classification models may desire to maximize the F1 score or MCC metrics, but the model with the lowest CE loss is not necessarily the one with the highest metrics. Moreover, a model can lower its CE loss by manipulating the predictions of samples that are already correctly separated. For an extreme example, consider a model that predicts 0.8 for all positive samples and 0.2 for all negative samples and a second model that predicts 0.99 for all positive samples and 0.01 for all negative samples. Both models perform perfect classification, yet the second model will have a lower CE loss. In a less extreme scenario, a model may improve its CE loss by simply manipulating the probabilities of easy-to-classify samples, but such manipulations would not improve real metrics such as the F1 score or MCC.
@@
There are multiple ways to remedy these issues. Lin et al. (2020) created the Focal Loss, which provides a correction factor that reduces the effect of increasing the certainty of a correct prediction, forcing the model to improve on difficult-to-classify samples to lower the focal loss. Berman et al. (2018) created the Lovász loss, which is an optimizable form of the Jaccardi Index and is claimed to perform better than the CE loss due to its better categorizing of small objects and lower false negatives. Other losses not evaluated in this work include the Baikal loss of Gonzalez & Miikkulainen (2020), the soft cross entropy of Ilievski & Feng (2017), the sigmoid F1 of Bénédict et al. (2022), the work of Lee et al. (2021), and the MCC loss of Abhishek & Hamarneh (2021). In particular, the approaches of this work resemble those of Lee et al. (2021) and Abhishek & Hamarneh (2021), but these references did not use weighting or focal transforms, and the definitions for the differentiable F1 and MCC loss functions used in these works are slightly different from the definitions in this work.
@@
The standard formulations of evaluation metrics (such as the F1 score or MCC) are not differentiable, so they cannot be directly used as loss functions. However, it is possible to modify them for differentiability. One method is treating the true / false categories as prediction probabilities instead of binary values. For example, if the model predicts for a positive (negative) sample, that sample would increase the true positives (negatives) by 0.7 and the false negatives (positives) by 0.3. This modification can be considered as a differentiable MCC loss function. The focal modification of Lin et al. (2020) can be added to this loss function by modifying to , which leads to the focal differentiable MCC loss function. Finally, it is possible to add class weighting: by multiplying the true positives and false negatives by a scalar W, it is possible to prioritize predictions on the positive class (W > 1) or negative class (W < 1). All the above changes combined lead to the weighted focal differentiable MCC loss function (Algorithm 1).
@@ 3.1 Transformer Encoders Have Good Predictive Power on This Dataset, but are Inferior to RNNs
The primary goal of this subsection is to compare transformer models with the best RNN from Seber & Braatz (2024) using the modified dataset of Wulff-Fuentes et al. (2021) (Section 2.1) and the weighted CE loss function. While our transformer models surpass models published before Seber & Braatz (2024) and achieve an F1 score equal to 24.31% and an MCC equal to 22.49% (Table 1), transformers exhibit inferior recall at the same precision level (Fig. 1) than the RNNs from Seber & Braatz (2024), leading to inferior F1 and MCC metrics. It should be noted that the work of Pokharel et al. (2023) used an ensemble of three transformer models to predict O-GlcNAcylation and obtained considerably inferior F1 and MCC metrics than the RNN of Seber & Braatz (2024) and the transformer models in this work. It is possible that transformers are an ill-suited architecture for this task, potentially due to the amount of data available, which is why we train primarily RNNs in the following sections.
Metric Transformer-15; 5 win (This Work) Transformer-15; 10 win (This Work) Transformer-15; 15 win (This Work) Best Threshold 0.60 0.20 0.50 Recall (%) 20.54 26.06 26.02 Precision (%) 15.63 13.77 21.28 F1 Score (%) 17.75 18.02 23.41 MCC (%) 15.50 16.08 21.36
A surprising result is how the performance of the transformer models does not change significantly with the hyperparameters. Using large hidden sizes or adding another linear layer after the encoders slightly degrades the performance, while increasing the number of attention heads (to 4–5) and the number of stacked encoder layers (to 4) marginally improves the model. The most surprising result is the lack of significant improvement with increasing window size. As in Seber & Braatz (2024), our models perform better with larger window sizes for sizes up to 20 (Fig. 1 and Table 1). The performance of a transformer encoder with a window size = 5 is only marginally inferior to the RNN from Seber & Braatz (2024) using the same window size. However, whereas the RNNs’ performances increase quite significantly with increasing window size, the performance of the transformers barely improves. The only significant performance increase comes from increasing the window size from 10 to 15, but barely any performance is gained from further increasing this window size to 20 (Fig. 1).